Two systems have the following equations of state, where R is the gas constant per mole. The mole numbers are N1 = 3 and N2 = 5. The initial temperatures are T1 = 175 K and T2 = 400 K. What is the temperature once equilibrium is reached (in Kelvin)?


1/T1=3/2×R×N1/U1

1/T2=5/2×R×N2/U2


1
Expert's answer
2020-09-07T08:18:03-0400

we know that

1/T=3/2R(n/U)

for two different temperature

1/T1=3/2R(n1/u1)

1/175=3/2 R(3/U1)

U1=6547.275

similarity for other system

1/T2=3/2R (N2/U2)

1/400=3/2 R(5/U2)

U2 =24942 J

so

U =U2_u1=24942 _6547.275

so during equilibrium condition

1/T=3/2R(n/U)

=3/2(8.314)(8/18394.725)

T = 184.375K